Calculate the pH after 0.020 mol of NaOH is added to 1.00 L of a solution consisting of 0.100 M HONH2 and 0.100 M HONH3Cl.
Answer #1
Kb for HONH2 = 1.1E-8
Ka for HONH3+ = Kw/Kb = 1E-14/1.1E-8 = 9.1E-7
pKa = -log Ka = -log (9.1E-7) = 6.04

If you add 0.02mol of OH-, your concentration of base will increase to 0.12M and your acid will decrease to 0.08M. Use the Henderson-Hasselbalch equationto find the pH:

pH = pKa + log (base/acid)
pH = 6.04 + log (0.12/0.08)
pH = 6.22
answered by: I Need Help